OpenStudy (anonymous):

limit lim x+2 dived square root x^2+5-(3) ->-2

OpenStudy (anonymous):

pls confirm, limit of (x+2)/sqrt(x^2+5 - 3)?

OpenStudy (anonymous):

lim x->-2

OpenStudy (anonymous):

but 3 is not under sgrt its out side

OpenStudy (anonymous):

is 3 included in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[= \lim_{x \rightarrow 2} \frac{x+2}{\sqrt{x^2+5} -3}\]

OpenStudy (anonymous):

apply L'Hospital's rule

OpenStudy (anonymous):

differentiate top and bottom as it gives 0/0 when you substitutue.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[= \lim_{x \rightarrow 2} \frac{1}{ \frac{x}{\sqrt{x^2+5}}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[= \frac{1}{ \frac{2}{3}} = \frac{3}{2}\]

OpenStudy (anonymous):

yes but negavit 3/2

OpenStudy (anonymous):

why

OpenStudy (anonymous):

because its x-> -2

OpenStudy (anonymous):

ohh yeh..

OpenStudy (anonymous):

thanks anyway

OpenStudy (anonymous):

bye